# GaussianOcc: Fully Self-supervised and Efficient 3D Occupancy Estimation with Gaussian Splatting

### v1

@@ Abstract
We introduce GaussianOcc, a systematic method that investigates the two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views. First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training. To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection. Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps, semantic maps), which is both time-consuming and less effective. We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting. As a result, the proposed GaussianOcc method enables fully self-supervised (no ground truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering). The relevant code will be available in https://github.com/GANWANSHUI/GaussianOcc.git.
Figure 1: Problem setting of GaussianOcc. Given the surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for the ground truth occupancy label and ground truth 6D pose for training.
@@ 1 Introduction
The existing methods [54, 24] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [13]. However, these approaches face two significant limitations. First, volume rendering is performed at real-world scale, which requires the availability of ground truth 6D pose to calculate the multi-view photometric loss across sequential images. Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [27, 1, 19], due to the dense sampling operation required. These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
To address the aforementioned limitations, we explore a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [27, 1]. Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network. This eliminates the need for a ground truth 6D pose during training. To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering. Instead, we propose performing Gaussian splatting directly from the 3D voxel space. In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we directly optimize the attributes of these Gaussians—such as semantic features and opacity—within the voxel space. Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
The main contributions of this work are summarized as follows.
  • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of the Gaussian splatting.
  • We propose Gaussian splatting for cross-view projection, which could provide the scale information to get rid of the need of a ground truth 6D pose during the training.
  • We propose Gaussian splatting from voxel space achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared with the volume rendering.
@@ 2.1 Surround view depth estimation
The surround view setting offers an ego-centric 360-degree perception solution [35, 39, 3, 14]. [17] introduces the surround view benchmark in the supervised setting, which learns the depth scale directly from the ground truth depth map. FSM [16] is a pioneering work in scale-aware surround view depth estimation relying on stereo constraint [53]. However, subsequent studies [44, 28] have found that reproducing the performance of [16] is challenging. Surrounddepth [44] improves the scale supervision signal using sparse point clouds from Structure-from-Motion (SFM). Building on the spatial and temporal constraints in FSM, [28] introduces a volume feature fusion module to enhance performance. For better performance, [38] proposes the temporal offline refinement strategy based on the multiple cameras and monocular depth refinement. These works [16, 44, 28, 38] all use traditional projection and index interpolation for cross-view synthesis, computing the loss between the synthesized view and target images. In contrast, our approach employs Gaussian splatting with a dedicated design, achieving better performance compared to the solely spatial constraint.
@@ 2.2 Surround view 3D occupancy estimation
Surround view 3D occupancy estimation has gained significant attention in recent years, with several benchmarks based on the Nuscene dataset [45, 41, 43]. In addition to the advanced architectures being proposed [32, 56, 51, 31], another research trend involves utilizing volume rendering for 3D supervision with 2D supervision [12, 37, 54, 24]. SimpleOcc [12] pioneered the use of volume rendering for 3D occupancy estimation, exploring both supervised and self-supervised learning. RenderOcc [37] extends semantic information for rendering. OccNeRF [54] and SelfOcc [24] share a similar approach by using 2D open-vocabulary semantic models to generate semantic maps for supervision. However, since volume rendering processes are conducted at real-world scale, these self-supervised methods [12, 54, 24] require ground truth 6D poses from sensors to provide the real-world scale for training. In contrast, we are exploring a solution that utilizes the overlap region in adjacent cameras to learn the real-world scale, eliminating the need for ground truth 6D poses.
@@
3D Gaussian splatting has become a popular method for modeling 3D and 4D scenes using well-posed images [27, 26, 8, 47], which has the property of fast rendering compared to the volume rendering in neural radiance field [36, 19, 11]. In the context of driving scenes, a line of research has focused on scene-specific reconstruction [22, 58, 49]. Our work, however, investigates the function of Gaussian splatting in a generalized setting. In this setting, existing methods generally construct 3D Gaussians from the unprojection of learned 2D Gaussian attributes [7, 57, 33, 40]. We also employ this unprojection approach, but uniquely, our approach constructs cross-view information from adjacent views to learn scale information through Gaussian splatting projection.
Additionally, two recent works, GaussianFormer [25] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction. However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
@@ 3 Method
Figure 2: The overview of the proposed GaussianOcc. With a sequence of surround images, we employ a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting. This approach provides scale information in the joint training with the 6D pose network (Stage 1). For 3D occupancy estimation, we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering (Stage 2). We omit the line to from the 6D pose net to the loss in stage 1 for clarity.
@@ 3.1 Preliminaries
In this paper, we investigate the use of 3D Gaussian splatting [27] for fully self-supervised 3D occupancy estimation. 3D Gaussian Splatting (3D-GS) is a technique for modeling static 3D scenes using point primitives. Each point is represented as a scaled Gaussian distribution characterized by a 3D covariance matrix $\Sigma$ and a mean $\mu$. The Gaussian distribution for a point $\mathcal{X}$ is given by:
$G(\mathcal{X})=e^{-\frac{1}{2}(\mathcal{X}-\mu)^{T}\Sigma^{-1}(\mathcal{X}-\mu)}$ (1)
To enable efficient optimization via gradient descent, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as follows:
$\Sigma=\mathbf{RSS}^{T}\mathbf{R}^{T}$ (2)
The projection of Gaussians from 3D space onto a 2D image plane involves a view transformation $\mathbf{W}$ and the Jacobian of the affine approximation of the projective transformation $\mathbf{J}$. The 2D covariance matrix $\Sigma^{\prime}$ is computed as:
$\Sigma^{\prime}=\mathbf{JW\Sigma W}^{T}\mathbf{J}^{T}$ (3)
Subsequently, an alpha-blend rendering technique, similar to that used in NeRF [36], is applied. This is formulated as:
$\mathbf{C}_{\text{color}}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i})$ (4)
Here, $c_{i}$ represents the color of each point, and the density $\alpha_{i}$ is determined by the product of a 2D Gaussian with covariance $\Sigma^{\prime}$ and a learned per-point opacity. The color is defined using spherical harmonics (SH) coefficients as described in [10, 27].
In summary, the basic 3D Gaussian Splatting approach characterizes each Gaussian point with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$.
3.2 Overview of the proposed GaussianOcc
The overview of the GaussianOcc is illustrated in Figure 2. The key contributions are the introduction of cross-view Gaussian splatting for scale-aware training and voxel grid Gaussian splatting for faster rendering. Unlike the original Gaussian splatting used for scene-specific 3D reconstruction, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids. This approach models the 3D scene in the 2D image plane as a depth map and in 3D grid space as a voxel (occupancy) format.
@@ 3.3 Scale-ware training by Gaussian Splatting
Scale from spatial camera rig: Similar to the previous work [16, 44], the scale information is from the surround camera rig. Specifically, the real-world scale could be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
@@
where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$th camera, $D_{t}^{i}$ is the predict depth map of $i$th camera, $p_{t}$ is the corresponding pixel during the warping. The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$. However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ is easy to be out of the image bounds with the weak supervision signal in such a small overlap region. Our practice also verifies this limitation. To provide a stronger supervision signal, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the scale-aware training, which is time-consuming and not straightforward. Differently, we propose to use the Gaussian splatting for scale-aware projection to achieve the cross-view stereo constraint.
Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we use a depth network to predict the Gaussian attributes in 2D grid space, including the depth map, scale map, and rotation map. For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions. Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section. We then perform splatting rendering on the adjacent views to obtain the rendered image. If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
The process of acquiring the overlap mask is shown in Figure 3. The acquisition of the overlap mask is based on volume rendering. We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view. Note that in the DDAD dataset [14], we exclude regions with self-occlusion (such as parts of the vehicle body). Finally, we apply an erosion operation from Open CV library [2] to the mask for purification.
@@ 3.4 Fast rendering by Gaussian Splatting
Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [12, 54, 24, 37] employed volume rendering based on dense sampling. While the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation (MLPs) [36], it remains time-consuming, particularly when incorporating semantic map rendering. For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066. However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,457,600$. This redundancy in densely sampled points aids optimization with volume rendering but is highly inefficient.
Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices. Therefore, we investigate the use of Gaussian splatting to further accelerate the rendering process. For each vertex in the 3D voxel grid, we obtain the necessary attributes for Gaussian splatting, such as 3D position, semantic information, and opacity. Consequently, we propose treating each vertex as a 3D Gaussian. Since these vertices are arranged in the 3D voxel space, we can model the 3D scene by optimizing the related Gaussian attributes.
@@ 3.5 Loss function
We formulate the the total loss function in:
$\mathcal{L}_{total}=\mathcal{L}_{temporal}+\mathcal{L}_{cross}+\mathcal{L}_{semantic}+\mathcal{L}_{tv},$ (6)
$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\beta\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(1-\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (7)
where the terms $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively. Note that $\hat{I}_{t}$ in the temporal-view photometric loss, $\mathcal{L}_{temporal}$, is generated by projecting pixels from the source image using the coordinate index. In contrast, $\hat{I}_{t}$ in the cross-view photometric loss, $\mathcal{L}_{cross}$, is derived from our proposed cross-view Gaussian splatting method. $\mathcal{L}_{semantic}$ is the 2D semantic loss. $\mathcal{L}_{tv}$ is the regularization loss for the whole voxel space. The balancing parameter $\beta$ is set to 0.85 to balance the two loss terms.

### v3

@@ Abstract
We introduce GaussianOcc, a systematic method that investigates two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views. First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training. To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection. Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps and semantic maps), which is time-consuming and less effective. We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting. As a result, the proposed GaussianOcc method enables fully self-supervised (no ground-truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering). The relevant code is available in https://github.com/GANWANSHUI/GaussianOcc.git.
Abstract
In this supplementary material, we provide more implementation details, experiment results, and further discussion on the limitations and future work.
@@ 1 Introduction
Figure 1: Problem setting of GaussianOcc. Given a surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for ground truth occupancy label and ground truth 6D pose for training.
@@
Existing methods [54, 23] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [12]. However, these approaches face two significant limitations. First, volume rendering is performed at real-world scale, which requires the availability of ground truth 6D pose to calculate the multi-view photometric loss across sequential images. Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [26, 1, 18], due to the dense sampling operation required. These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
To address the aforementioned limitations, we propose a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [26, 1]. Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network. This eliminates the need for ground truth 6D pose during training. To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering. Instead, we propose performing Gaussian splatting directly from the 3D voxel space. In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we optimize the attributes of these Gaussians—such as semantic features and opacity—directly within the voxel space. Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
In summary, our core contributions are as follows:
  • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of Gaussian splatting.
  • We propose Gaussian splatting for cross-view projection module, which can provide scale information to get rid of the need of ground truth 6D pose during the training.
  • We propose Gaussian splatting from voxel space module, achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared to the previous works with volume rendering.
@@ 2.1 Surround view depth estimation
The surround view setting offers an ego-centric 360-degree perception solution [34, 39, 3, 13]. [16] introduces the surround view benchmark in the supervised setting, which learns the depth scale directly from the ground truth depth map. FSM [15] is a pioneering work in scale-aware surround view depth estimation relying on stereo constraint [53]. However, subsequent studies [44, 27] have found that reproducing the performance of [15] is challenging. Surrounddepth [44] improves the scale supervision signal using sparse point clouds from Structure-from-Motion (SFM). Building on the spatial and temporal constraints in FSM, [27] introduces a volume feature fusion module to enhance performance. For better performance, [38] proposes the temporal offline refinement strategy based on the multiple cameras and monocular depth refinement. These works [15, 44, 27, 38] all use traditional projection and index interpolation for cross-view synthesis, computing the loss between the synthesized view and target images. Compared to the traditional projection, our approach employs Gaussian splatting with a dedicated design for cross-view constraint, achieving better performance.
@@ 2.2 Surround view 3D occupancy estimation
Surround view 3D occupancy estimation has gained significant attention in recent years, with several benchmarks based on the nuScenes dataset [45, 41, 43]. In addition to the advanced architectures being proposed [31, 56, 51, 30], another research trend involves utilizing volume rendering for 3D occupancy learning with 2D supervision [11, 37, 54, 23]. SimpleOcc [11] pioneered the use of volume rendering for 3D occupancy estimation, exploring both supervised and self-supervised learning. RenderOcc [37] extends semantic information for rendering. OccNeRF [54] and SelfOcc [23] share a similar approach by using 2D open-vocabulary semantic models to generate semantic maps for supervision. However, since volume rendering processes are conducted at real-world scale, these self-supervised methods [11, 54, 23] require ground truth 6D poses from sensors to provide the real-world scale for training. Differently, we are exploring a solution that utilizes the overlap region in adjacent cameras to learn the real-world scale, eliminating the need for ground truth 6D poses.
Figure 2: GaussianOcc is a two-stage method. In Stage 1, we train a scale-aware 6D pose network, using a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting. This approach provides scale information in the joint training with the 6D pose network. Based on the 6D pose from Stage 1, we perform self-supervised 3D occupancy estimation in Stage 2, where we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering. Note that, for clarity, we omit the line from the 6D pose network to the loss in Stage 1, and the 2D encoder is independent for each stage (not shared).
@@
3D Gaussian splatting has become a popular method for modeling 3D and 4D scenes using well-posed images [26, 25, 8, 47], which has the property of fast rendering compared to the volume rendering in neural radiance field [35, 18, 10]. In driving scenes, a line of research has focused on scene-specific reconstruction [21, 58, 49]. Our work, however, investigates the function of Gaussian splatting in a generalized setting, where existing methods generally construct 3D Gaussians from the unprojection of learned 2D Gaussian attributes [7, 57, 32, 40]. We also employ this unprojection approach, but uniquely, our approach constructs cross-view information from adjacent views to learn scale information through Gaussian splatting projection. In addition, we investigate the Gaussian splatting on the voxel space for faster rendering compared with the previous volume rendering-based works [11, 54, 23].
Note that, two recent works, GaussianFormer [24] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction. However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
@@ 3.1 Preliminaries
3D Gaussian splatting [26] is for modeling static 3D scenes using point primitives, where each primitive is characterized with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$. The original Gaussian splatting is for scene-specific, fast 3D novel view synthesis, where the attributes of Gaussian points are optimized by the multi-view constraint. In contrast, in this work, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids, allowing us to benefit from the Gaussian splatting rendering for the scale-aware training by cross-view constraint and faster rendering on voxel grids as illustrated in Figure 2.
@@ 3.2 Scale-ware training by Gaussian Splatting
Scale from spatial camera rig: Similar to the previous work [15, 44, 27], the scale information is from the surround camera rig. Specifically, the real-world scale can be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
@@
where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$-th camera, $D_{t}^{i}$ is the predicted depth map of $i$-th camera, $p_{t}$ is the corresponding pixel during the warping. The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$. However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ can easily go out of image bounds with the weak supervision signal in such a small overlap region. Therefore, apart from the spatial loss, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the direct depth supervision to provide a stronger supervision signal, but it is time-consuming and not straightforward. Different from [44], [27] enhances the depth estimation performance with spatio-temporal context that does not need the sparse depth from SFM but the performance is still limited. Inspired by the explicit sparse depth supervision in [44], we ask whether we can enforce the cross-view constraint on adjacent views more explicitly. The answer is yes. We find that the nature of Gaussian splatting is scale-aware projection that could serve for the cross-view stereo constraint. We propose Gaussian splatting for projection in stage 1 for better scale-aware training.
Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we adapt a depth network [52] to predict the Gaussian attributes in 2D grid space, where, apart from the original depth map, we also predict the scale map and rotation map. For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions. Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section (Table 3). We then perform splatting rendering on the adjacent views to obtain the rendered image. If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
The process of acquiring the overlap mask is illustrated in Figure 3. The acquisition of the overlap mask is similar to volume rendering. We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view. The overlap mask is determined by the camera extrinsic and defined max depth (e.g., 80m in nuScenes). Note that in the DDAD dataset [13], we exclude regions with self-occlusion (such as parts of the vehicle body). Finally, we apply an erosion operation from OpenCV library [2] to the mask for purification.
@@ 3.3 Fast rendering by Gaussian Splatting
Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [11, 54, 23, 37] employed volume rendering based on dense sampling.Although the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation [35], it remains time-consuming, particularly when incorporating semantic map rendering. For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066. However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,577,525$. This redundancy in densely sampled points helps optimization with volume rendering but is highly inefficient.
Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices. Interestingly, we find that it is suitable to use the Gaussian splatting to replace the volume rendering if we regard the vertices of each the voxel grid as the position of the 3D Gaussian. Then, we can optimize the attributes of the 3D Gaussian, such as semantic and opacity information. For example, in the empty space of the voxel, even though we have the vertices at that region during the splatting rendering, after the optimization, the network would predict the opacity as zero at these vertices then these vertices would not contribute any geometry or semantic information during the rendering. Since all the vertices are arranged in 3D voxel space with real-world 3D position $\mathcal{X}$, we use fixed scale $s$ and fixed rotation $r$ for each vertex for simplification. This allows us to model the 3D scene by optimizing the rest Gaussian attributes (semantic and opacity).
@@ 3.4 Loss function
We formulate the loss function of each stage as follows:
$\mathcal{L}_{stage1}=\mathcal{L}_{temporal}+\mathcal{L}_{cross},$ (2)
$\mathcal{L}_{stage2}=\mathcal{L}_{temporal}+\mathcal{L}_{semantic},$ (3)
$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (4)
where $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively. Note that $\hat{I}_{t}$ in the temporal-view photometric loss $\mathcal{L}_{temporal}$ is generated by projecting pixels from the source image using the coordinate index. In contrast, $\hat{I}_{t}$ in the cross-view photometric loss $\mathcal{L}_{cross}$ is derived from our proposed cross-view Gaussian splatting method. $\mathcal{L}_{semantic}$ is 2D semantic loss. The $\beta$ is set to 0.15 to balance the two loss terms. For $\mathcal{L}_{temporal}$ and $\mathcal{L}_{semantic}$, we recommend [54] for further details.